# SOLD: Reinforcement Learning with Slot Object-Centric Latent Dynamics

### v1

@@ Abstract
Learning a latent dynamics model provides a task-agnostic representation of an agent’s understanding of its environment. Leveraging this knowledge for model-based reinforcement learning holds the potential to improve sample efficiency over model-free methods by learning inside imagined rollouts. Furthermore, because the latent space serves as input to behavior models, the informative representations learned by the world model facilitate efficient learning of desired skills. Most existing methods rely on holistic representations of the environment’s state. In contrast, humans reason about objects and their interactions, forecasting how actions will affect specific parts of their surroundings. Inspired by this, we propose Slot-Attention for Object-centric Latent Dynamics (SOLD), a novel algorithm that learns object-centric dynamics models in an unsupervised manner from pixel inputs. We demonstrate that the structured latent space not only improves model interpretability but also provides a valuable input space for behavior models to reason over. Our results show that SOLD outperforms DreamerV3, a state-of-the-art model-based RL algorithm, across a range of benchmark robotic environments that evaluate for both relational reasoning and low-level manipulation capabilities.
Videos are available at https://slot-latent-dynamics.github.io/.
@@ 1 Introduction
Advances in reinforcement learning (RL) have showcased the ability to learn sophisticated control strategies through interaction, achieving superhuman performance in domains ranging from board games (Silver et al., 2016) to drone racing (Kaufmann et al., 2023). While these approaches excel in settings where explicit models of the environment are available or abundant data can be collected, learning complex control tasks in a sample-efficient manner remains a significant challenge. Model-based RL (MBRL) has emerged as a promising approach to address this limitation by constructing models of the environment’s dynamics. Notably, the Dreamer framework (Hafner et al., 2020; 2021; 2023) has demonstrated improved sample efficiency over model-free baselines by learning behaviors solely through imagined rollouts.
@@
While these research efforts have produced world models capable of accurately predicting the dynamics of visual tasks, they rely on a holistic representation of the environment’s state. In contrast, humans perceive the world by parsing scenes into individual objects (Spelke, 1990), anticipating how their actions will influence specific components of their surroundings. This ability is crucial in complex tasks that require reasoning about multiple objects and their interactions, which is common in robotic manipulation. Learning structured representations of an environment through objects introduces a powerful inductive bias. The learned representations not only enhance the interpretability of the dynamics prediction but also foster decision-making by providing a structured input space for behavior models to reason over. Despite these advantages, the integration of object-centric representations and world models remains largely underexplored. To the best of our knowledge, no prior work has introduced a method that performs object-centric model-based RL directly from pixels.
@@
To address the limitations of holistic representations in model-based RL, we propose Slot-Attention for Object-centric Latent Dynamics (SOLD), a novel algorithm that leverages structured, object-centric states within the latent space of its world model. Our method introduces two key innovations to the model-based RL framework. The first contribution is a dynamics model that predicts future frames in terms of their slot representation. To achieve this, we extend OCVP (Villar-Corrales et al., 2023) into an action-conditional model by projecting action commands into the slot dimension. A transformer-based (Vaswani et al., 2017) transition model processes the set of object-slots and actions to predict the subsequent frame. Notably, the dynamics model is trained solely from pixels through a loss on the reconstructions and slot representations of the predicted frames. The second key component is a transformer-based model backbone, which we call Slot Aggregation Transformer, that is used to make time-step-wise predictions from the history of object slots. Specifically, it serves as the architectural backbone to the reward, value, and action model, allowing us to perform model-based RL training on the basis of object-centric representations.
@@
For systematic evaluation, we introduce a suite of visual robotics tasks, shown in Figure 1, that require varying levels of relational reasoning and manipulation capabilities. We perform an extensive comparison on this benchmark, demonstrating that our method achieves superior performance to the state-of-the-art MBRL algorithm DreamerV3 (Hafner et al., 2023). Furthermore, we apply SOLD to tasks from two RL benchmarks that were not designed to be object-centric, providing evidence of the generalizability of our framework. In summary, we make the following contributions:
  • We introduce SOLD, which is to the best of our knowledge, the first object-centric model-based RL algorithm that learns entirely from pixel inputs.
  • Our method outperforms DreamerV3 across a range of visual robotics environments, excelling in tasks that require both relational reasoning and low-level manipulation skills.
@@
  • We overcome limitations of prior object-centric methods in RL. Namely, we show that our object-centric encoder-decoder module can be adapted to state distributions vastly different from those seen under a random policy during pre-training – an essential capability for solving many complex tasks. Additionally, we show its generalization potential to environments not explicitly designed for object-centric reasoning.
@@
Figure 1: We evaluate SOLD on diverse visual environments from (left to right) our multi-object robotic control benchmark, Meta-World (Yu et al., 2019), and DM-Control (Tassa et al., 2018).
2 Background
@@
SOLD employs SAVi (Kipf et al., 2022), an encoder-decoder architecture with a structured bottleneck composed of $N$ permutation-equivariant object embeddings denoted as slots, in order to recursively parse a sequence of video frames ${\bm{o}}_{0:\tau}=\{{\bm{o}}_{0},...,{\bm{o}}_{\tau}\}$ into their object representations ${\bm{Z}}_{0:\tau}=\{{\bm{Z}}_{0},...,{\bm{Z}}_{\tau}\},{\bm{Z}}_{t}\in\mathbb{R}^{N\times D_{Z}}$. At time $t$, SAVi encodes the input video frame ${\bm{o}}_{t}$ into a set of feature maps ${\bm{F}}_{t}\in\mathbb{R}^{L\times D_{h}}$, where $L$ is the size of the flattened grid (i.e. $L=\text{width}\cdot\text{height}$), and uses Slot Attention (Locatello et al., 2020) to iteratively refine the previous slot representations conditioned on the current features. Slot Attention performs cross-attention between the slots and image features with the attention coefficients normalized over the slot dimension, thus encouraging the slots to compete to represent feature locations:
@@
where $q$ and $k$ are learned linear mappings to a common dimension $D$. The slots are then independently updated via a shared Gated Recurrent Unit (Cho, 2014) (GRU) followed by a residual MLP:
${\bm{Z}}_{t}\doteq\text{MLP}(\text{GRU}({\bm{A}}\cdot v({\bm{F}}_{t}),{\bm{Z}}_{t-1}))\;\;\text{with}\;{\bm{A}}_{n,l}\doteq\frac{{\bm{A}}_{n,l}}{\sum_{i=0}^{L-1}{\bm{A}}_{n,i}},$ (2)
where $v$ is a learned linear projection. The steps described in Equations 1 and 2 can be repeated multiple times with shared weights to iteratively refine the slots and obtain an accurate object-centric representation of the scene. Finally, SAVi independently decodes each slot of ${\bm{Z}}_{t}$ into per-object images and alpha masks, which can be normalized and combined via weighted sum to render video frames. SAVi is trained end-to-end in a self-supervised manner with an image reconstruction loss.
OCVP
Our dynamics model builds on OCVP (Villar-Corrales et al., 2023) in order to autoregressively predict future object slots conditioned on past object states. OCVP is a transformer-encoder model that decouples the processing of object dynamics and interactions, thus leading to interpretable and temporally consistent object predictions while retaining the inherent permutation-equivariant property of the object slots. This is achieved through the use of two specialized self-attention variants: temporal attention updates a slot representation by aggregating information from the corresponding slot up to the current time step, without modeling interactions between distinct objects, whereas relational attention models object interactions by jointly processing all slots from the same time step.
(a) World Model Learning: SAVi encodes images ${\bm{o}}_{t}$ into slots ${\bm{Z}}_{t}$, which are predicted by the dynamics model given history of slots and actions. We reconstruct the images and compute their actual slot representation to shape the dynamics prediction.
(b) Behavior Learning: The actor and critic are trained via imagined rollouts in the latent space of the world model. Trajectories start after $S$ seed frames (visualized for $S=1$) and predict forward with actions sampled from the actor network.
@@ 3 Slot-Attention for Object-centric Latent Dynamics
We propose Slot-Attention for Object-centric Latent Dynamics (SOLD), a method that combines model-based RL akin to the Dreamer framework (Hafner et al., 2020; 2021; 2023) with object-centric representations. The three core components of our algorithm are: the object-centric world model, which predicts the effects of actions on the environment, the critic, which estimates the value of a given state, and the actor that selects actions to maximize this value.
@@ 3.1 World Model Learning
Most prior works rely on generating a single, holistic representation of the environment’s state, which contrasts with findings from cognitive psychology. Humans perceive scenes as compositions of objects (Spelke, 1990) and reason about how their actions affect distinct parts of their environment. Furthermore, environment dynamics can be compactly explained in terms of objects and their interactions Battaglia et al. (2016). Therefore, we propose to structure the latent space by decomposing visual environments into their constituent parts.
@@
To create a world model that operates on object-centric latent representations, we build on top of OCVP (Villar-Corrales et al., 2023). We begin by pretraining SAVi on a dataset of $10^{6}$ frames from random episodes. Having a sufficiently large initial dataset is crucial for meaningful object-centric representations to emerge. We do not freeze the pretrained encoder-decoder models, allowing slots to adapt to novel configurations that do not occur during random pre-training. The sequence of object slots ${\bm{Z}}_{0:t}$ alongside the action commands ${\bm{a}}_{0:t}$ serve as inputs to our transformer-based dynamics model which predicts the slot representation of the next frame $\hat{{\bm{Z}}}_{t+1}$:
$\makebox[120.00018pt][l]{Encoder:}&&&{\bm{Z}}_{t}=e_{\eta}({\bm{o}}_{t}),\\ \makebox[120.00018pt][l]{Decoder:}&&&\hat{{\bm{o}}}_{t}=d_{\eta}({\bm{Z}}_{t}),\\ \makebox[120.00018pt][l]{Dynamics model:}&&&\hat{{\bm{Z}}}_{t+1}=p_{\psi}({\bm{Z}}_{0:t},{\bm{a}}_{0:t}),\textrm{and}\\ \makebox[120.00018pt][l]{Reward predictor:}&&&\hat{r}_{t}\sim p_{\zeta}(\hat{r}_{t}\mid{\bm{Z}}_{0:t}).\end{aligned}$ (3)
@@
For the dynamics model, we follow the sequential attention pattern proposed in Villar-Corrales et al. (2023), which disentangles relational and temporal attention to decouple the processing object dynamics and interactions. During training, we provide the slot representation of $S$ seed frames as context. We append the predictions to the context and apply this process in an autoregressive manner to predict the subsequent $T$ frames. We do not employ teacher forcing so that the dynamics model learns to handle its own imperfect predictions. To shape the predicted representations, we reconstruct the subsequent frame $\hat{{\bm{o}}}_{t+1}$ and extract the SAVi representations of the actual frame ${\bm{Z}}_{t+1}$ to compute the hybrid dynamics loss:
@@
For all loss terms, we specify the parameter group that is being optimized and omit stop-gradients for other models to avoid cluttering the notation.
@@
The reward predictor solves a regression problem where the prediction depends on the slot representations but is not directly tied to any single slot. To address this, we introduce the Slot Aggregation Transformer (SAT) as an architectural backbone, which introduces output tokens and a variable number of register tokens for all time-steps. Register tokens, recently shown to enhance computation in vision transformers (Darcet et al., 2024), can aid computation when processing a set of inputs to produce a singular output. To encode the position information, we adopt ALiBi (Press et al., 2022) in place of absolute position encoding. ALiBi introduces linear biases directly into the attention scores, effectively encoding token recency. This approach helps to generalize to sequences longer than those seen during training. A detailed description of the SAT can be found in Section B.3. To efficiently represent a wide range of reward values, we avoid directly predicting a scalar reward. Instead, the MLP head $f_{\zeta}$ outputs logits of a softmax distribution over $K$ exponentially spaced bins $b_{i}$. The predicted reward can then be computed as the expectation over these bins:
$\bm{b}\doteq\operatorname*{symexp}([-20,...,+20]),\hskip 30.00005pt\hat{r}_{t}\doteq\operatorname*{softmax}(f^{\text{MLP}}_{\zeta}({\bm{h}}_{t}))^{T}\bm{b},$ (5)
@@ 3.2 Behavior Learning
$R_{t}^{\lambda}\doteq\hat{r}_{t+1}+\gamma\left((1-\gamma)\hat{R}_{t+1}+\lambda R_{t+1}^{\lambda}\right)\hskip 5.69046pt\text{where}\hskip 5.69046ptR_{T}^{\lambda}\doteq\hat{R}_{T},$ (8)
which is trained to minimize the resulting loss:
@@
The actor is optimized to select actions that maximize its expected return while encouraging exploration through an entropy regularizer. Its model architecture is similar to the critic and reward predictor, but instead of regressing a scalar value, it predicts the parameters of the action distribution. Specifically, the MLP head outputs the mean $\bm{\mu}_{t}$ and standard deviation $\bm{\sigma}_{t}$ to parameterize a normal distribution $\mathcal{N}(\bm{\mu}_{t},\bm{\sigma}_{t}|{\bm{Z}}_{0:t})$ over possible actions. The trade-off in the actor’s loss function weights expected returns with maintaining randomness in the actor outputs and is hence subject to reward scale and frequency of the current environment. To adapt to varying scales of value estimates across different environments, we use a normalization factor $\text{scale}_{V}$:
$\mathcal{L}_{\text{actor}}(\theta)\doteq-\sum_{t=0}^{T-1}\frac{\hat{R}_{t}^{\lambda}}{\max(1,\text{scale}_{V})}+\eta\mathrm{H}(\mathcal{N}(\bm{\mu}_{t},\bm{\sigma}_{t}\mid{\bm{Z}}_{0:t})),$ (10)
@@
$\text{scale}_{V}\doteq\mathrm{EMA}\left(\mathrm{Per}(\hat{R}_{t}^{\lambda},95)-\mathrm{Per}(\hat{R}_{t}^{\lambda},5),0.99\right).$ (11)

### v2

# SOLD: Slot Object-Centric Latent Dynamics Models for Relational Manipulation Learning from Pixels

@@ Abstract
Learning a latent dynamics model provides a task-agnostic representation of an agent’s understanding of its environment. Leveraging this knowledge for model-based reinforcement learning (RL) holds the potential to improve sample efficiency over model-free methods by learning from imagined rollouts. Furthermore, because the latent space serves as input to behavior models, the informative representations learned by the world model facilitate efficient learning of desired skills. Most existing methods rely on holistic representations of the environment’s state. In contrast, humans reason about objects and their interactions, predicting how actions will affect specific parts of their surroundings. Inspired by this, we propose Slot-Attention for Object-centric Latent Dynamics (SOLD), a novel model-based RL algorithm that learns object-centric dynamics models in an unsupervised manner from pixel inputs. We demonstrate that the structured latent space not only improves model interpretability but also provides a valuable input space for behavior models to reason over. Our results show that SOLD outperforms DreamerV3 and TD-MPC2 – state-of-the-art model-based RL algorithms – across a range of benchmark robotic environments that require relational reasoning and manipulation capabilities. Videos are available at https://slot-latent-dynamics.github.io/.
Machine Learning, ICML
@@ 1 Introduction
Advances in reinforcement learning (RL) have showcased the ability to learn sophisticated control strategies through interaction, achieving superhuman performance in domains ranging from board games (Silver et al., 2016) to drone racing (Kaufmann et al., 2023). While these approaches excel in settings where explicit models of the environment are available or abundant data can be collected, learning complex control tasks in a sample-efficient manner remains a significant challenge. Model-based RL (MBRL) has emerged as a promising approach to address this limitation by constructing models of the environment dynamics. Notably, the Dreamer framework (Hafner et al., 2019, 2020, 2023) has demonstrated improved sample efficiency over model-free baselines by learning behaviors solely through imagined rollouts.
@@
While these research efforts have produced world models capable of accurately predicting the dynamics of visual tasks, they rely on a holistic representation of the environment state. In contrast, humans perceive the world by parsing scenes into individual objects (Spelke, 1990), anticipating how their actions will influence specific components of their surroundings. Relational reasoning, particularly in environments with multiple interacting objects, is a cornerstone of human intelligence and a crucial capability for robotic manipulation tasks (Battaglia et al., 2018). Introducing structured, object-centric representations into MBRL represents a powerful inductive bias, enabling agents to reason about task-relevant components of the environment while ignoring irrelevant details. Such structured representations not only enhance interpretability but also improve the efficiency of behavior learning by providing models with meaningful latent spaces. Despite these advantages, the integration of object-centric representations and world models remains largely underexplored. To the best of our knowledge, no prior work has introduced a method that performs object-centric model-based RL directly from pixels.
@@
To address the limitations of holistic representations in MBRL, we propose SOLD, a novel algorithm that leverages structured, object-centric states within the latent space of its world model. The contributions of our method are twofold. First, we introduce an object-centric dynamics model that predicts future frames in terms of their slot representation. Building on the OCVP framework (Villar-Corrales et al., 2023), we introduce an action-conditional dynamics model, enabling accurate forecasting of action effects on individual objects. Notably, the dynamics model is trained solely from pixels through a loss on the reconstructions and slot representations of the predicted frames, bypassing the need for object annotations and facilitating scalability to complex visual tasks. Second, we propose the Slot Aggregation Transformer, a novel architectural backbone that aggregates information from the history of object slots to make reward, value, and action predictions. This enables efficient MBRL training grounded in structured, interpretable representations.
@@
For systematic evaluation, we introduce a suite of visual robotics tasks, shown in Figure 1, that require varying levels of relational reasoning and manipulation capabilities. We perform an extensive comparison on this benchmark, demonstrating that our method achieves superior performance to both state-of-the-art MBRL algorithms DreamerV3 (Hafner et al., 2023) and TD-MPC2 (Hansen et al., 2024). Further, to highlight its broader potential, we apply SOLD to tasks from two RL benchmarks that are not object-centric by design, providing evidence of the generalizability of our framework. In summary, we make the following contributions:
  • We introduce SOLD, the first object-centric MBRL algorithm to learn directly from pixel inputs, achieving state-of-the-art performance on visual robotics tasks that require both relational reasoning and manipulation.
  • By visualizing learned attention weights, we show that our method produces human-interpretable attention patterns, providing insights into the decision-making process of behavior models.
@@
  • We overcome limitations of prior object-centric RL methods by showing that our encoder-decoder module can (i) be adapted to state distributions vastly different from those seen under random pre-training, and (ii) generalize to environments that are not object-centric by design.
2 Background
@@
Figure 1: Suite of visual environments requiring relational reasoning and low-level manipulation to be solved.
@@
SOLD employs SAVi (Kipf et al., 2022), an encoder-decoder architecture with a structured bottleneck composed of $N$ permutation-equivariant object embeddings, referred to as slots. It recursively parses a sequence of video frames ${\bm{o}}_{0:\tau}={{\bm{o}}_{0},...,{\bm{o}}_{\tau}}$ into their object representations ${\bm{Z}}_{0:\tau}={{\bm{Z}}_{0},...,{\bm{Z}}_{\tau}},{\bm{Z}}_{t}\in\mathbb{R}^{N\times D_{Z}}$. At time $t$, SAVi encodes the input video frame ${\bm{o}}_{t}$ into a set of feature maps ${\bm{F}}_{t}\in\mathbb{R}^{L\times D_{h}}$, where $L$ is the size of the flattened grid (i.e. $L=\text{width}\cdot\text{height}$), and uses Slot Attention (Locatello et al., 2020) to iteratively refine the previous slot representations conditioned on the current features. Slot Attention performs cross-attention between the slots and image features with the attention coefficients normalized over the slot dimension, thus encouraging the slots to compete to represent feature locations:
@@
where $q(.)$ and $k(.)$ are learned linear mappings to a common dimension $D$. The slots are then independently updated via a shared Gated Recurrent Unit (Cho, 2014) (GRU) followed by a residual MLP:
$\begin{split}{\bm{Z}}_{t}&\doteq\text{MLP}(\text{GRU}({\bm{A}}\cdot v({\bm{F}}_{t}),{\bm{Z}}_{t-1})),\\ {\bm{A}}_{n,l}&\doteq\frac{{\bm{A}}_{n,l}}{\sum_{i=0}^{L-1}{\bm{A}}_{n,i}},\end{split}$ (2)
and $v(.)$ is a learned linear projection. The steps described in Equations 1 and 2 can be repeated multiple times with shared weights to iteratively refine the slots and obtain an accurate object-centric representation of the scene.
Finally, SAVi independently decodes each slot of ${\bm{Z}}_{t}$ into per-object images and alpha masks, which can be normalized and combined via weighted sum to render video frames. SAVi is trained end-to-end in a self-supervised manner with an image reconstruction loss.
(a) World Model Learning: SAVi encodes images ${\bm{o}}_{t}$ into slots ${\bm{Z}}_{t}$, which are predicted by the dynamics model given the history of slots and actions. We reconstruct the images and compute their actual slot representation to shape the dynamics prediction.
(b) Behavior Learning: Actor and critic are trained via imagined rollouts in the latent space of the world model. Trajectories start after $S$ seed frames (visualized for $S=1$) and predict forward with actions ${\bm{a}}_{t}$ sampled from the actor network.
@@
Object-Centric Video Prediction (OCVP)
Our dynamics model builds on OCVP (Villar-Corrales et al., 2023) in order to autoregressively predict future object slots conditioned on past object states. OCVP is a transformer-encoder model (Vaswani et al., 2017) that decouples the processing of object dynamics and interactions, thus leading to interpretable and temporally consistent object predictions while retaining the inherent permutation-equivariant property of the object slots. This is achieved through the use of two specialized self-attention variants: temporal attention updates a slot representation by aggregating information from the corresponding slot up to the current time step, without modeling interactions between distinct objects, whereas relational attention models object interactions by jointly processing all slots from the same time step.
Figure 3: Open-loop predictions of our object-centric dynamics model. Starting from a single context frame, our model predicts the next 50 frames by propagating slot representations forward without access to any intermediate images.
@@ 3 Slot-Attention for Object-centric Latent Dynamics
We propose SOLD, a method that combines model-based RL with object-centric representations. The three core components of our method are: the object-centric world model, which predicts the effects of actions on the environment, the critic, which estimates the value of a given state, and the actor that selects actions to maximize this value.
@@ 3.1 World Model Learning
Most prior works rely on generating a single, holistic representation of the environment state, which contrasts with findings from cognitive psychology. Humans perceive scenes as compositions of objects (Spelke, 1990) and reason about how their actions affect distinct parts of their environment. Furthermore, environment dynamics can be compactly explained in terms of objects and their interactions (Battaglia et al., 2016). Therefore, we propose to structure the latent space by decomposing visual environments into their constituent parts.
@@
To create a world model that operates on object-centric latent representations, we build on top of OCVP (Villar-Corrales et al., 2023). We begin by pretraining SAVi on a static dataset of frames collected from random episodes. Having a sufficiently large initial dataset is crucial for meaningful object-centric representations to emerge. These pretrained representations serve as the foundation for SOLD’s object-centric world model. However, we do not freeze the pretrained encoder-decoder models, allowing slots to adapt to novel configurations that do not occur during random pre-training. The sequence of object slots ${\bm{Z}}_{0:t}$ alongside the action commands ${\bm{a}}_{0:t}$ serve as inputs to our transformer-based dynamics model which predicts the slot representation of the next frame $\hat{{\bm{Z}}}_{t+1}$:
$\makebox[80.00012pt][l]{Encoder:}&&&{\bm{Z}}_{t}=e_{\eta}({\bm{o}}_{t}),\\ \makebox[80.00012pt][l]{Decoder:}&&&\hat{{\bm{o}}}_{t}=d_{\eta}({\bm{Z}}_{t}),\\ \makebox[80.00012pt][l]{Dynamics model:}&&&\hat{{\bm{Z}}}_{t+1}=p_{\psi}({\bm{Z}}_{0:t},{\bm{a}}_{0:t}),\;\textrm{and}\\ \makebox[80.00012pt][l]{Reward predictor:}&&&\hat{r}_{t}\sim p_{\zeta}(\hat{r}_{t}\mid{\bm{Z}}_{0:t}).\end{aligned}$ (3)
@@
For the dynamics model, we use the sequential attention pattern proposed by Villar-Corrales et al. (2023), which disentangles relational and temporal attention to decouple object dynamics and interactions. During training, we provide the slot representation of $S$ seed frames as context. We append the predictions to the context and apply this process in an autoregressive manner to predict the subsequent $T$ frames. We do not employ teacher forcing so that the dynamics model learns to handle its own imperfect predictions. To shape the predicted representations, we reconstruct the subsequent frame $\hat{{\bm{o}}}_{t+1}$ and extract the SAVi representations of the actual frame ${\bm{Z}}_{t+1}$ to compute the hybrid dynamics loss:
@@
For all loss terms, we specify the parameter group that is being optimized and omit stop-gradient notations for other models to avoid cluttering the notation.
@@
The reward predictor solves a regression problem that maps slot representations to scalar reward values, where the prediction depends on the set of slots rather than being tied to any specific one. To address this, we introduce the Slot Aggregation Transformer (SAT) as an architectural backbone, which introduces output tokens and a variable number of register tokens for all time-steps. Register tokens, recently shown to enhance computation in vision transformers (Darcet et al., 2024), can aid computation when processing a set of inputs to produce a singular output. To encode positional information, we adopt ALiBi (Press et al., 2022) in place of absolute position encoding. ALiBi introduces linear biases directly into the attention scores, effectively encoding token recency. This approach helps to generalize to sequences longer than those seen during training. A detailed description of the SAT can be found in Section D.3. To efficiently represent a wide range of reward values, we avoid directly predicting a scalar reward. Instead, the MLP head $f_{\zeta}$ outputs logits of a softmax distribution over $K$ exponentially spaced bins $b_{i}$. The predicted reward can then be computed as the expectation over these bins:
$\begin{split}\bm{b}&\doteq\operatorname*{symexp}([-20,...,+20]),\\ \hat{r}_{t}&\doteq\operatorname*{softmax}(f^{\text{MLP}}_{\zeta}({\bm{h}}_{t}))^{T}\bm{b},\end{split}$ (5)
@@ 3.2 Behavior Learning
$R_{t}^{\lambda}\doteq\hat{r}_{t+1}+\gamma\left((1-\lambda)\hat{R}_{t+1}+\lambda R_{t+1}^{\lambda}\right),$ (8)
where $R_{T}^{\lambda}\doteq\hat{R}_{T}$, which is trained to minimize the resulting loss:
@@
The actor is optimized to select actions that maximize its expected return while encouraging exploration through an entropy regularizer. Its model architecture is similar to the critic and reward predictor, but instead of regressing a scalar value, it predicts the parameters of the action distribution. Specifically, the MLP head outputs the mean $\bm{\mu}_{t}$ and standard deviation $\bm{\sigma}_{t}$ of a normal distribution $\mathcal{N}(\bm{\mu}_{t},\bm{\sigma}_{t}|{\bm{Z}}_{0:t})$ over possible actions. The trade-off in the actor’s loss function weights expected returns with maintaining randomness in the outputs and is hence subject to reward scale and frequency of the current environment. To adapt to varying scales of value estimates across different environments, we use a normalization factor $\text{s}_{V}$:
$\mathcal{L}_{\text{actor}}(\theta)\doteq-\sum_{t=0}^{T-1}\frac{\hat{R}_{t}^{\lambda}}{\max(1,\text{s}_{V})}+\eta\mathrm{H}(\pi_{\theta}(\bm{a}_{t}\mid{\bm{Z}}_{0:t})),$ (10)
@@
$\text{s}_{V}\doteq\mathrm{EMA}\left(\mathrm{Per}(\hat{R}_{t}^{\lambda},95)-\mathrm{Per}(\hat{R}_{t}^{\lambda},5),0.99\right).$ (11)